George Ferris built the first Ferris wheel for the World’s Exposition in Chicago in 1892. Its diameter was 250 ft. If this wheel rotated once every 5 minutes, what was the speed of a car on the circumference of the wheel in feet per second?
A) 157 ft/min B) 15.7 ft/sec C) 2.618 ft/min D) 2.618 ft/sec E) 157 min/hr

Answer by Theo(2978)

diameter is 250 feet.
radius is half that = 125 feet.
circumference is 2*pi*r = 2*pi*125 = 785.3981634 feet = 1 rotation of the ferris wheel.
785.3981634 / 300 = 2.617993878 feet per second.

that would be selection D.